Calculate the weight of a 68.0 kg person:

a) on the surface of the earth

b) at the top of Everest (8848 m above sea level).

c) has 2 1/2 times the radius of the earth


1
Expert's answer
2021-07-15T10:51:19-0400

(a)

"W=mg=68.0*9.81=667\:\rm N"

(b)

"W=mg\frac{R^2}{(R+H)^2}=68.0*9.81\frac{6378^2}{(6378+8.848)^2}=665\:\rm N"

(c)

"W=mg\frac{R^2}{(R+H)^2}=68.0*9.81\frac{R^2}{(R+2.5R)^2}=54.5\:\rm N"